OpenStudy (fredfredburgeryes):

is this graph symmetric about the axis? r=9sin7theta not sure how this goes, can someone walk me through this?

OpenStudy (fredfredburgeryes):

is it symmetric about x, y and/or origin?

OpenStudy (priyar):

Do you know how the sine graph is?

OpenStudy (michele_laino):

hint: we have this: \[\begin{gathered} \sin \left( {7\theta } \right) = 7{\left( {\cos \theta } \right)^6}\left( {\sin \theta } \right) - 35{\left( {\cos \theta } \right)^5}{\left( {\sin \theta } \right)^2} + 21{\left( {\cos \theta } \right)^4}{\left( {\sin \theta } \right)^3} - \hfill \\ - 21{\left( {\cos \theta } \right)^2}{\left( {\sin \theta } \right)^5} + {\left( {\sin \theta } \right)^7} \hfill \\ \end{gathered} \] wherein: \[\begin{gathered} x = r\cos \theta \hfill \\ y = r\sin \theta \hfill \\ \end{gathered} \] so we can rewrite the function as follows: \[{r^8} = 9{r^7}\sin \left( {7\theta } \right)\] and finally: \[{\left( {{x^2} + {y^2}} \right)^4} = 9\left\{ {7{x^6}y - 35{x^5}{y^2} + 21{x^4}{y^3} - 21{x^2}{y^5} + {y^7}} \right\}\]

OpenStudy (fredfredburgeryes):

oh my, thanks!